### v1

# supernovae: general — supernovae: individual (SNe 2022esa, 2018ibb, 2022jli) — stars: evolution — binaries: general

Abstract
A class of supernovae (SNe) termed ‘SN Ic-CSM’ are characterized by late-time emergence of narrow emission lines of elements formed in the oxygen core of a massive star. A popular scenario is the interaction of the SN ejecta and O-rich circumstellar medium (CSM), i.e., Circumstellar Interaction (CSI). Uncovering the progenitor system of SNe Ic-CSM plays a critical role in understanding the final evolution of a massive star to a bare C+O star. In this Letter, we present observations of SN 2022esa which we show is an SN Ic-CSM. Surprisingly, a stable periodicity of days is found in its light-curve evolution with a hint of a slowly increasing period over days. We argue that the main power source is likely the interaction of the SN ejecta and O-rich CSM, while the energy input by the post-SN eccentric binary interaction within the SN ejecta is another possibility. In either case, we propose a massive Wolf-Rayet (WR) star as the progenitor, in a WR-WR or WR-BH (black hole) binary that will eventually evolve to a BH-BH binary. Specifically, in the CSI scenario, the progenitor system is an eccentric binary system with an orbital period of about a year, leading to the observed periodicity through the modulation in the CSM density structure. We also show that some other objects, superluminous SN I 2018ibb (a pair-instability SN candidate) and peculiar SN Ic 2022jli (the first example showing stable periodic modulation), show observational similarities to SNe Ic-CSM and may be categorized as SN Ic-CSM variants. Complemented with a large diversity in their light-curve evolution, we propose that SNe Ic-CSM (potentially linked to SNe Ibn/Icn) are a mixture of multiple channels that cover a range of properties in the progenitor star, the binary companion, and the binary orbit.
1 Introduction
Recent development in transient observations has revealed a rich diversity of a transient zoo (e.g., [schulze2025]). ‘SN Ic-CSM’ is a recently found class of objects, characterized by late-time emergence of narrow ( a few 1,000 km s-1) emission lines of elements formed in the oxygen core of a massive star (e.g., O and Mg) (kuncarayakti2022)111SN Ib 2014C (margutti2017) and SN Ic 2017dio (kuncarayakti2018) showed transition to type IIn with strong Balmer emissions in the late-phase. In this paper, they are not classified as SN Ib-CSM or Ic-CSM, which we define are characterized by narrow emission lines in the H-poor environment.. Within a still very limited sample, the best studied cases, SNe 2021ocs and 2022xxf, both showed a transition from type Ic (an explosion of a C+O star) to SN Ic-CSM (kuncarayakti2023), establishing their association with C+O star progenitors. In addition to the spectral evolution, they are distinct in the light-curve (LC) evolution from SNe Ic. The interaction between the SN ejecta and CSM (Circumstellar Interaction: CSI), both rich in oxygen, has been proposed as a power source.
The progenitor channel toward SNe Ic-CSM has not been clarified; it is not even clear whether SNe Ic-CSM form a single population. Also, a relation between SNe Ic-CSM and Ibn/Icn (intense interaction with H-deficient CSM only in the time scale of weeks; pastorello2007; gal-yam2022) has not been clarified.
SN 2022esa was discovered on 2022 March 12 (MJD 59650.61)222The date is shown in UT unless mentioned. by Asteroid Terrestrial-impact Last Alert System (ATLAS) (tonry2022), in 2MFGC 13525 at . It was classified as SN Ia-CSM (H-rich CSI-powered SN Ia; dilday2012) by lu2022. In this Letter, we present follow-up observations of SN 2022esa. Its spectral evolution reveals that SN 2022esa is indeed a new member of the SN Ic-CSM class333Recently, griffith2025 also mentioned that it is a peculiar SN Ic.. Its LC also shows unique features, including periodic modulation. We present these characteristic observational features, and discuss its main power source and the progenitor system.
2 Observations and Data Reduction
Spectra of SN 2022esa were obtained on 2022 May 28 ( d444In this Letter, we define the phase () of SN 2022esa as measured from the discovery date. The timescale (e.g., LCs shown in figures) is shown in the SN rest frame, corrected for redshift, for all the objects.) and June 2 ( d), using Kyoto Okayama Optical Low-dispersion Spectrograph with optical-fiber Integral Field Unit (KOOLS-IFU; matsubayashi2019) mounted on the Kyoto University 3.8m Seimei Telescope (kurita2020). We used the VPH-blue grism with the wavelength coverage of Å and the spectral resolution of . The total exposure time was 2,700 s in each night. A late-time spectrum was obtained on 2023 June 13 ( d), using Faint Object Camera And Spectrograph (FOCAS) mounted on the Subaru telescope. We used a 0.8" slit and the B300 grism with no order-cut filter, covering 3650–8300Å with the spectral resolution of . The total exposure time was 3,600 s. Additionally, we obtained (120 s) and -band (180 s) images on the same night with FOCAS.
The reduction of the FOCAS data was performed in the standard manner with IRAF555IRAF is distributed by the National Optical Astronomy Observatory, which is operated by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation. (see maeda2022_Iax). The reduction of the KOOLS-IFU data follows similar procedures, with the Hydra package (hydra) and custom routines666http://www.o.kwasan.kyoto-u.ac.jp/inst/p-kools/reduction-201806/index.html.
The spectra of SN 2022esa (figure 1) show a number of narrow emission lines from an underlying unresolved HII region. For the FOCAS spectrum, the contamination is especially substantial given the decreasing SN flux, and thus we subtracted it with a spectrum of a nearby HII region taken in the same 2D image. The flux was converted to the absolute scale by assuming the distance modulus of 35.24 mag and the Milky-Way (MW) extinction with mag. We assume that the extinction within the host is negligible; in addition to the absence of strong Na ID absorption, the Balmer decrement measured from the underlying HII region is consistent with the MW-only extinction.
The light curves (LCs) of SN 2022esa (figure 2) were constructed as follows. For the early phase up to days, we used public resources. The ATLAS and -band magnitudes were obtained using ATLAS forced photometry server777https://fallingstar-data.com/forcedphot/ (tonry2018; shingles2021). The ZTF and -band magnitudes were obtained through the ALeRCE explorer888https://alerce.science/ (foster2021). We used the FOCAS spectrum to estimate the ‘SN-only’ magnitude in the late phase ( d). First, spectrophotometry on the ‘SN + HII region’ resulted in and mag (in Vega), being consistent with the photometry performed on the and images ( and mag). We then performed spectrophotometry on the HII region-subtracted spectrum, obtaining the pure-SN magnitudes of and mag (in AB).
We also constructed infrared LCs using the data from the NEOWISE mission (mainzer2014) in the W1 (3.4 µm) and W2 (4.6 µm) bands, taken by the Wide-field Infrared Survey Explorer (WISE; wright2010). We performed point-spread function (PSF) photometry at the position of the source in the NEOWISE difference images (de2020; de2023).
3 Results
3.1 Modulation in the Light Curve
Figure 2 shows bumpy structures in the LC evolution, the characteristic that is rarely seen in SNe. The evolution is coherent in different bands; excluding the Atlas -band data that are not well sampled, we performed the periodgram analysis. We first fit the overall LC of a given band by a polynomial function of 5th order, creating the mean LC. This mean LC was subtracted from the original LC, and then the residual was further divided by the mean LC. The resulting ‘fractional-residual’ LCs for the atlas -band and ZTF - and -bands are shown in figure 3 (left). Surprisingly, SN 2022esa showed clear periodicity over multiple cycles, which has been reported previously only for peculiar SN Ic 2022jli (moore2023; chen2024)999Another example, type I superluminous SN (SLSN-I) 2024afav, has recently been reported to show a quasi-periodic LC (farah2025)..
We performed the Lomb-Scargle (LS) periodogram analyses for the residual LCs separately in the -, -, and -band data, using LombScargle module implemented in astropy. This was done iteratively by rejecting outliers by 3 clipping in the data-model residual. A sharp peak in the power spectrum is found at the period that is consistent among all the three bands; days (), days (), and days () (figure 3, middle). To further investigate possible quasi-periodic behavior, we performed the sliding-window analysis with a window size of 100 days and a step size of 25 days on the ZTF- data. As shown in figure 3 (right), the data are consistent with no change in the period within 1, while there is a hint of a possible increase in the period over time101010The sliding-window analysis of the Atlas- data resulted in the same conclusion, with a larger error in the derived period..
In case of SN 2022jli, the origin of the modulation has not been conclusively clarified (moore2023; cartier2024), but the ‘post-SN’ binary interaction between the compact object left behind the explosion and the bound companion star has been proposed (chen2024). As compared to SN 2022jli, the modulation seen here has several distinct properties. SN 2022jli showed (1) a shorter period (12.4 d), and (2) an asymmetric temporal profile with a fast rise ( d) while it is consistent with a symmetric profile for SN 2022esa. Further, (3) the modulation was observed in the late phase after d in SN 2022jli, while the early-rising LC of SN 2022esa is consistent with having the modulation already. There are also other differences between SNe 2022esa and 2022jli in the overall LC and spectral evolutions, which will be discussed in subsequent sections.
Temporal variations are sometimes seen in radio LCs of stripped-envelope SNe (SESNe, i.e., SNe IIb/Ib/Ic) (wellon2012), and the quasi-periodic modulation has been reported for a few SNe (ryder2004; kotak2006; soderberg2006). The behavior has been generally interpreted to be caused by the modulation in the CSM density structure. Especially popular is the scenario that involves a luminous blue variable-like progenitor (kotak2006). Indeed, the temporal variation in the LC, while usually without periodic behavior, is common among SNe Ic-CSM (see section 3.3); the H-poor CSI in a non-smooth CSM distribution is a popular scenario (kuncarayakti2023).
3.2 SN 2022esa as a new member of SNe Ic-CSM
SN 2022esa showed broad spectral features already in the early phase with no clear sign of H from the SN (figure 1). The broad features were further developed at the maximum phase. Finally, in the late epoch, strong and narrow (but resolved) lines (FWHM km s-1) from intermediate mass elements (IMEs) such as O and Ca were clearly detected; these properties are inconsistent with SN Ia-CSM, but rather indicate that SN 2022esa was associated with O-rich environments, i.e., SN Ic or its variants.
We compare the spectral evolution of SN 2022esa with those of SNe Ic-CSM and a few related objects (figure 4). SN Icn 2021ckj is a rapidly decaying SN where the interaction with ‘confined’ C-rich CSM is the main energy source (pellegrino2022; nagao2023). SNe 2010mb (ben-ami2014), 2021ocs (kuncarayakti2022), and 2022xxf (kuncarayakti2023) were classified as SNe Ic-CSM based on their late-time spectra. They showed flat or slowly decaying LCs even with multiple peaks. From these features, SNe Ic-CSM have been proposed to be SNe Ic exploded within dense and extended O-rich CSM. SNe 2021ocs and 2022xxf did not show the signatures of strong CSI in the early phase, and were classified as canonical SN Ic and broad-lined SN Ic (SN Ic-BL) initially, pointing to the diverse nature of the SN ejecta in the SN Ic-CSM class. SN 2019tsf is a peculiar SN Ib in its LC evolution (sollerman2020), later showing the spectral features of SNe Ic-CSM (pyykkinen2025). SLSN-I 2018ibb has been suggested to be a pair-instability SN (PISN) (schulze2024; nagele2024), with the late-time spectrum indicating an additional power provided by the H-poor CSI (chugai2024). Finally, SN Ib 2022jli, the first SN showing the clear periodicity, is also shown for comparison (moore2023; chen2024).
SN 2022esa shows striking similarity to SNe Ic-CSM. The earliest spectrum is overall similar to the spectrum of SN 2022xxf at the first peak. The broad features also show resemblance to SN Icn 2021ckj, while the narrow emission lines from highly-ionized carbon are missing in SN 2022esa. These similarities indicate that SN 2022esa is related to an oxygen-star progenitor, i.e., ‘SN Ic’.
The spectrum around the peak of SN 2022esa shows clear similarity to those of SNe Ic-CSM 2021ocs and 2022xxf. Interestingly, the candidate PISN 2018ibb and the ‘periodic’ SN Ic 2022jli are similar to SN Ic-CSM objects and SN 2022esa, inferring a link of these peculiar objects to the SN Ic-CSM class; we suggest that they could indeed be variants of SNe Ic-CSM, potentially sharing a similar power source and/or spectral-formation process.
[…]
3.3 Possible power sources
The large luminosity ( mag) and long duration ( mag for days) indicate the radiated energy of SN 2022esa reaching at least erg. In addition, the color remained rather constant and blue ( and mag) throughout the period up to d. The CSI is a promising candidate to explain these properties. Figure 5 compares the LCs of SN 2022esa and related objects, including ‘H-rich CSI’ SN IIn 2010jl (zhang2012) and Ia-CSM 2020uem (uno2023). In addition to the lack of the H features in the spectra, the LC evolution also indicates that SN 2020esa is neither of the SN Ia-CSM nor SN IIn class111111We however note that the LCs of SNe IIn are very diverse..
Some SNe Ic-CSM show two or more peaks, which has been associated with two different energy sources – 56Ni/Co/Fe heating at the first peak, and the O-rich CSI thereafter. The most likely energy source for SN 2022esa is thus the O-rich CSI, and the LC periodicity reflects the fluctuation in the CSM density. Under this scenario, we have two additional constraints; (1) the large peak luminosity indicates a very high mass-loss rate, reaching yr-1 (maeda2013_10jl; uno2023) just before the explosion (unlike SNe 2021ocs and 2022xxf showing the late-time emergence of the CSI), and (2) it requires a stable periodic phenomenon to introduce the CSM density fluctuation (unlike SN 2022xxf showing a bumpy LC but without clear periodicity).
We note two additional hints supporting the CSI scenario: (1) The properties of the radio emission (griffith2025), erg s-1 Hz-1 at yrs at 6 GHz, is typical of SNe IIn (chevalier2006; maeda2013_radio). (2) IR excess is seen in the late phase (figure 2), similar to those frequently observed in SNe IIn. The IR emission signals dusty environments associated with dense CSM either through an echo from pre-existing CS dust (at a larger distance; maeda2015_echo) or through the reprocessing of the optical photons to the IR by newly formed dust (at a smaller scale; maeda2013_10jl). Given the blue optical color even in the late phase, the pre-existing dust is a more likely scenario. We will further investigate the nature of the IR excess in a separate paper.
If the pre-SN mass loss would be perfectly periodic with no change in the mass-loss velocity (), the CSM ‘shells’ are equally spaced. Then, the observed LC modulation period is inversely proportional to the SN shock velocity . Therefore, we expect that the period is constant (for constant ) or increasing (for decreasing ). For example, for the decelerated SN shock in a steady-state wind-like CSM, where is the density slope in the SN ejecta (). Figure 3 (right) shows the predictions for and covering a range of SN-ejecta properties (chevalier1982). The observed period is consistent with this picture (i.e., no change within the error, but with a hint of an increasing period).
The ‘post’-SN binary interaction scenario proposed for SN 2022jli, where the ‘central engine’ through the post-SN binary interaction powers the LC periodically, is not rejected. However, the main argument against the CSI scenario for SN 2022jli, namely the variability shorter than the light-travel time, does not apply for SN 2022esa. Also, given its high luminosity, SN 2022esa would require a very strong binary interaction and a high mass-transfer rate, and the period may likely decrease rapidly, inconsistent with the constant or slowly increasing period in SN 2022esa.
3.4 SN-site environment
We measured metallicity at the SN site and the nearby H II region, using the O3N2 and N2 calibrations from marino13. We found that the metallicities for the two regions are the same within an error. Two calibrations provide consistent results; dex (N2) or dex (O3N2), i.e., . This is similar to the cases of some strongly interacting SNe, e.g. Type IIn SN 2010jl (stoll11; maeda2013_10jl).
The equivalent width of the H emission line (HEW) is frequently used to infer the age of the stellar population and the lifetime of the progenitor star (see e.g. hk18-ifu). The measured HEW of SN 2022esa environment is around 600Å. Using the Starburst99 single-burst model (leitherer99) at half solar metallicity, this corresponds to a stellar population age of around 4.7 Myr, thus to the lifetime of a M⊙ star. Such a massive star is expected to evolve to a massive WR star (heger2003). While the single-burst model can overestimate the progenitor mass, the young stellar population age suggests a massive star origin for SN 2022esa.
4 Discussion
4.1 Progenitor System: WR-WR or WR-BH binary?
Under the CSI scenario, the LC covering yr reflects the activity of the progenitor (system) in the final yrs yrs, where km s-1 and km s-1 are assumed. The LC modulation period is translated into yr for the variability period of the pre-SN mass loss. A binary interaction, where one component is the C+O-star SN progenitor, is an appealing possibility to create such a stable periodicity.
The circularized close binary, which is believed to be a major channel toward SESNe (fang2019), would not create a stable periodicity. Alternatively, a highly eccentric binary experiences repeated binary interactions every time the system passes the peri center, potentially resulting in a periodic mass loss. This suggests that the progenitor of SN 2022esa likely evolved to a C+O star without strong binary interaction (which circularizes the orbit) – a massive WR star that has lost the outer H and He envelopes by its strong wind. Assuming the total mass of the pre-SN binary systems as , the average pre-SN binary separation must have been . This again points to a massive WR star progenitor that can avoid becoming a red supergiant, thus not experiencing a Roche-lobe overflow and/or common envelope (CE) that will circularize the orbit and lead to a compact binary.
No trace of hydrogen may suggest an H-deficient companion star, i.e., either a WR-WR system (analogous to eccentric WR-WR binary Apep; callingham2020) or a WR-BH system, noting that a neutron-star (NS) companion is unlikely if the progenitor is a massive WR. While the (still rare) observed WR-BH systems generally show close circularized orbits (prestwich2007; binder2021), it may simply be due to observational biases; for O star-WR binaries (as a progenitor to WR-BH binaries) such as WR140, a period of a year to decades with an eccentric orbit is common (e.g., williams2009). The ejecta mass of SN 2022esa would not exceed half of the total mass for the binary system considered here, and the binary likely survived after SN 2022esa. Therefore, the proposed WR-WR or WR-BH system will ultimately lead to the formation of a BH-BH binary.
The LC decay of SN 2022esa is much faster than SNe interacting with extended CSM like SNe 2010jl and 2020uem. This indicates a steep CSM density gradient and an increasing mass-loss rate toward the SN. The binary interaction thus became ‘on’ in the last decade before the SN. Such a synchronization requires that the final evolution of the progenitor drove the binary interaction. One possibility is inflation of the progenitor star or some sort of activity in the final phase (ouchi2019) that could lead to an increasingly strong binary interaction toward the SN. For example, a massive WR progenitor has been proposed for SNe Ibn/Icn (pastorello2007; gal-yam2022) – putting such a star in an eccentric binary would lead to further enhancement of the mass loss and could lead to objects like SN 2022esa.
In the ‘post-SN’ binary interaction scenario, the observed period reflects the post-SN binary period – the average separation of . Given the lack of H, the companion star would also be a C+O star; it is unlikely that both stars evolve to C+O stars through strong binary interaction, and a massive WR-WR binary is a favored system also in this scenario. A few caveats should be noted; the emergence of H might be a transient feature and could simply be missed, or the H may be hidden as the binary interaction happens deep inside the ejecta. In these cases, a WR-O/B star binary progenitor system would also be possible. Either way, again the system may ultimately evolve to a BH-BH binary.
4.2 SN Ic-CSM diversities: Multiple populations?
[…]
SNe Ic-CSM shows a huge diversity in the LC evolution (figure 5); the peak magnitude, the time scale, and the evolution morphology are diverse. This may indicate that SNe Ic-CSM are likely a mixture of different populations originated in multiple progenitor channels. For most of them, the O-rich CSI is likely the main power source, with diverse properties in the CSM (and thus in the mass-loss history). Even the underlying power source may not be unique, possibly including the post-SN binary interaction. These channels may include a massive WR-WR or WR-BH binary (proposed here for SN 2022esa), a standard (relatively low-mass) SESN binary channel but extreme conditions for the binary mass transfer (perhaps involving a CE), a WR + normal (perhaps O) star binary (proposed for SN 2022jli), and a PISN (proposed for 2018ibb). In most cases, the binarity seems to be a key; depending on the nature of the binary components (types of stars and masses) and the orbit (separation and eccentricity), a huge diversity within the ‘SN Ic-CSM’ class (and possibly including Ibn/Icn) may be created.
5 Concluding Remarks
In this Letter, we presented unique features of SN 2022esa. In addition to the bright optical emission, it showed a stable periodicity ( d) in the LC, with a possible hint of a slowly increasing period over days. Further, the spectra evolved into the characteristic ‘SN Ic-CSM’ late-time spectrum, showing blue Fe bumps and narrow emission lines from IMEs.
For the power source, both the CSI and the post-SN binary interactions remain possible, while the CSI is more likely. In either case, we suggest a massive WR progenitor star, in a binary with either another WR or a BH. For the CSI scenario, we further require a highly eccentric orbit. Therefore, SN 2022esa may represent a progenitor system that will eventually evolve to a BH-BH binary.
We showed that the SNe Ic-CSM class has a huge diversity, pointing out that SNe 2018ibb (a PISN candidate) and 2022jli (another example showing a clear periodicity) have similarities to SNe Ic-CSM. We suggest that SNe Ic-CSM represent a mixture of multiple populations. The progenitor mass may span from the ‘canonical SESN progenitor’ () to a massive WR star (), or even to the PISN regime (). Furthermore, for many of them, a binary interaction seems to play a key role, adding a source of the diversity depending on the binary configuration that may further connect them to SNe Ibn/Icn.
SNe Ic-CSM can thus represent a rich and diverse zoo of massive binary evolution, including progenitors toward BH-BH binaries. As such, further studying them, both in observationally and theoretically, will provide a major contribution to uncover the still-unclarified final evolution of massive stars, roles of binarity, and formation channels of double compact-object binaries.
The authors thank Kentaro Aoki for his excellent assistance for the observation with the Subaru telescope (under the program S23A-023), and Niko Pyykkinen for providing the unpublished spectrum of SN 2019tsf. The data from the Seimei telescope were obtained under the KASTOR (Kanata And Seimei Transient Observation Regime) project (under the programs 22A-N-CT09 and 22A-K-0030). The Seimei telescope is jointly operated by Kyoto University and the Astronomical Observatory of Japan (NAOJ), with assistance provided by the Optical and Near-Infrared Astronomy Inter-University Cooperation Program. Some public data are obtained from WISeREP (https://www.wiserep.org).
Funding
K.M. acknowledges support from JSPS KAKENHI grant (JP24KK0070, JP24H01810, JP24K00682, and JP20H00174). H.K. was funded by the Academy of Finland projects 324504 and 340613. The work is partly supported by the JSPS Open Partnership Bilateral Joint Research Projects between Japan and Finland (K.M and H.K; JPJSBP120229923).
Data availability
The data underlying this article will be available through WiseRep.